# An angular-momentum preserving dissipative model for the point-mass $N$-body problem

### 12 Mar 2026 v1

@@ 4.1 Poincaré compactification
Figure 3: Phase space portraits for the compactified system (33) for $C=\gamma=1$. (A) The system without dissipation ($\alpha=0$). Dashed lines show the stable and unstable manifolds for $(w_{v\infty},z_{v\infty})=(0,0)$ and solid lines a few illustrative orbits. (B) The system with dissipation ($\alpha=0.5$). The equilibrium point $(w_{e},z_{e})=(0,0.5)$ divides the phase space, as indicaded by its manifolds (dotted pink lines).
Finally, there is another chart ($V_{2}$ in [Dumortier_book], p. 152) that contains the limit $v\to$-$\infty$ with r=r=\,constant, which we denote $v_{-\infty}$. This point is also an equilibrium with the same local dynamics as that near $(w_{v\infty},z_{v\infty}$). The vector field on this new chart is obtained by means of the change of varibles $(w,z)\mapsto(-w,-z)$ in vector field (33) and multiplied by 1-1, due to time reparametrization (last paragraph in pg. 152 of [Dumortier_book]).
Figure 4: Sketches of the phase space topology on the Poincaré disk for the Kepler system. Green dots mark the circular equilibrium $(r_{c},v_{c})$. (A) Conservative case ($\alpha=0$) showing the parabolic separatrix (red). (B) Dissipative case ($\alpha=0.1$) showing the separatrix and two center manifolds. (C) Dissipative case ($\alpha=0.5$) showing a new ejection center manifold and the connection of the center manifold from $(r_{e},v_{e})$ to $(r_{c},v_{c})$.
@@ 4.1.1 Phase portrait topology
To build the full phase portrait topology on the Poincaré disk, we considered three scenarios: $\alpha=0$, $\alpha=0.1$, and $\alpha=0.5$, all for $C=\gamma=1$ (see Fig. 4). The same topological results may be obtained for any other values of $C,\gamma>0$.
For $\alpha=0$, the global portrait is assembled from panels (A) of figures 1, 2, and 3. Figure 4 (A) shows the flow lines on the Poincaré half-disk, with a center at the circular orbit $(r_{c},v_{c})$, surrounded by periodic orbits (ellipses), together with parabolic and hyperbolic orbits that are homoclinic to the degenerate equilibrium at $r_{\infty}$. In this case the characteristic curves $\Gamma_{u(s)}$ from $(y_{\infty},x_{\infty})$ merge (red curve), forming the separatrix between bounded and unbounded orbits.
For $\alpha=0.5$, the global portrait is assembled from panels (B) of figures 1, 2, and 3. Figure 4 (C) shows the half-disk flow, where now a branch leaving $(r_{e},v_{e})$ spirals toward the circular orbit $(r_{c},v_{c})$, which is now an attractor. A second branch connects scattering (escape) orbits that are not captured by the attractor.
For $\alpha=0.1$, the phase portraits resemble those for $\alpha=0.5$, except that Fig. 2 (B) is replaced by Fig. 5, where a transition half-way from the conservative to $\alpha=0.5$ takes place. Here the separatrix is still present and coexists with the weak-center-manifold branch connecting $r_{\infty}$ to the circular attractor. The difference from the case $\alpha=0.5$ is that now three different regions appear in the phase space, as the weak-center-manifold from $v_{e}$ connects to $r_{\infty}$, separating captured and scattered orbits. When $\alpha=0.5$, this connection is lost, and only two regions appear, as the branch leaving $v_{e}$ connects to the circular attractor. The global portrait for this case in shown in Fig. 4 (B).
The threshold between these two scenarios occur for some $\alpha\in(0.1,0.5)$, where the center-manifold of $v_{e}$ merges to the characteristic curve $\Gamma_{s}$ of $r_{\infty}$. At this point the scattering region disappears, with all incoming orbits ($v<0$) being captured and lead to circularization. The basin of attraction of $r_{c}$ increases along with $\alpha$, as expected.
Figure 5: Phase space portrait for the compactified system (30) for $C=\gamma=1$ and $\alpha=0.1$, showing that the weak-center-manifold, associated with $(r_{e},v_{e})=(0,2)$, which in $(y,x)$ is a point at infinity, is located within the $S_{a}$ sector and connects to $(y_{\infty},x_{\infty})=(0,0)$, instead of $(y_{c},x_{c})=(1,0)$ as seen for $\alpha=0.5$ (Fig. 2 (B) and (C)).
@@ 4.2 The angular-momentum-energy space
An orbit in the conservative Kepler system maps to a single point in the CH space, as $C$ and $\mathcal{H}$ are constants. In the presence of the dissipation considered here, however, the image of an initial point, not on the curve of minimum energy, always moves vertically down a line. There are two possibilities for this motion as $t\to\infty$. One, where the point moves vertically down till it reaches the minimum energy curve, which happens if, and only if, the initial point is in the basis of attraction of the circular orbit. The second possibility is that the point moves vertically down until reaching a limit with $\mathcal{H}>0$. This scenario happens if, and only if, the initial point belongs either to the scattering region or to the escaping region.
Smale [Smale] provides a thorough analysis on this direction for the conservative case of systems with $N$-bodies.
@@ 6 Conclusion
In the derivation of these equations, if one is able to show that the central dissipative force can be expressed as a function of the gravitational potential for a generic degree of homogeneity, the techniques and results obtained here immediately follow for any value of $d$.
@@
As mentioned above, the Kepler system corresponds to the homographic dynamics of any planar central configuration under dissipation. For sufficiently small dissipation and nonzero angular momentum, any such configurations in the circular equilibrium basins of attraction will have their homothetic (radial) motion damped. This leads any central configuration to purely rotational motion and therefore to a relative equilibrium. We showed that the circularization dynamics inside these basins are governed by a weak-center-manifold.
With the addition of bodies in the system ($N>2$), the lack of a minimum to the energy changes the dynamics in the circularization attractor basin, as dissipation goes on forever and collisions may occur. Despite the rapidly growing complexity with the addition of bodies, even in the conservative case, the simpler two-bodies scenario may hint on the stabilization of orbits near large primaries for long periods of time.
To extend the current model, one could consider the effect of rotation of bodies, reintroducing some of the neglected tidal effects. Moreover, to increase flexibility, one may also assing different dissipation strengths amongst pairs of bodies, $\alpha\to\alpha_{ij}=\alpha_{i}\alpha_{j}$, adapting to realistic scenarios where the primary bodies respond differently to tides due to its inner structure.

### 06 Apr 2026 v2

@@ 4.1 Poincaré compactification
Figure 3: Phase space portraits for the compactified system (36) for $C=\gamma=1$. (A) The system without dissipation ($\alpha=0$). Dashed lines show the stable and unstable manifolds for $(w_{v\infty},z_{v\infty})=(0,0)$ (blue point) and solid lines a few illustrative orbits. (B) The system with dissipation ($\alpha=0.5$). The equilibrium $(w_{e},z_{e})=(0,0.5)$ (red point) divides the phase space, as indicated by its manifolds (dotted pink lines).
Finally, we analyze the limit $v\to$-$\infty$ that, after the change of variables in equation (35), corresponds to the quadrant $w\leq 0$, $z\leq 0$ in equation (36). We denote the equilibrium at $v=$-$\infty$ as
$(w_{v-\infty},z_{v-\infty})=(0,0)$
A desingularization of $(w_{v-\infty},z_{v-\infty})$ is achieved using variables similar to those in equation (38)
$(w,z)\longmapsto(p,q)\quad\text{where}\quad w=pq^{3},\;\;z=-q^{2}$ (40)
We are interested in the quadrant $p\leq 0$, $q\geq 0$.
In the new variables, the vector field in equation (36) becomes, after division by $q^{6}/2$,
[…]
This equation has a single equilibrium $(p,q)=(0,0)$, which is a saddle. Its unstable manifold lies on the $p$-axis, and its stable manifold lies on the $q$-axis. As expected, the outcome of this analysis is the phase space portrait shown in the quadrant $w\leq 0$, $z\leq 0$ in Fig. 3  [^2] An information provided by the blow up is that in the quadrant $w\leq 0$, $z\leq 0$ solutions initially close to the equilibrium $(0,0)$ approach it approximately along straight lines $z(t)\approx\frac{z(0)}{w(0)}w(t)$ before diverging along the $w$-axis. This can be observed in Fig. 3..
Figure 4: Sketches of the flow line topology on the Poincaré disk for the Kepler system. Green dots mark the circular equilibrium $(r_{c},v_{c})$. (A) Conservative case ($\alpha=0$ or $c=\infty$), showing the parabolic separatrix (pink), which corresponds to the merging of the stable $\Gamma_{s}$ and unstable $\Gamma_{u}$ curves of the “equilibrium” $r_{\infty}$, see equation (4.1). (B) Weakly dissipative case ($\alpha=0.1$ or $c\approx 2.15$), showing that $\Gamma_{s}$ (pink) is homoclinic to $r_{\infty}$, while $\Gamma_{u}$ (green) is attracted to $(r_{c},v_{c})$. All points inside $\Gamma_{s}$ converge to the circular orbit as $t\to\infty$ and diverge to $r=\infty$ as $t\to$-$\infty$. The center-unstable manifold $W^{c}_{e}$ (red) of the ejection point $v_{e}$ converges to $r_{\infty}$ as $t\to\infty$. Orbits outside $\Gamma_{s}$ and below $W^{c}_{e}$ belong to the scattering region, being bi-asymptotic to $r_{\infty}$ as $t\to\pm\infty$. Orbits above $W^{c}_{e}$ form the escaping region, being asymptotic to $r_{\infty}$ as $t\to\infty$ and to $v_{\infty}$, as $t\to$-$\infty$, see equation (37). (C) Strongly dissipative case ($\alpha=0.5$ or $c\approx 1.26$), showing that $\Gamma_{s}$ (pink) is asymptotic to $v_{\infty}$ as $t\to$-$\infty$. Orbits above $\Gamma_{s}$ belong to the escaping region, while those below $\Gamma_{s}$ lie in the basin of attraction of the circular orbit. The center-unstable manifold $W^{c}_{e}$ (red) is attracted to the circular orbit as well as $\Gamma_{u}$ (green). Orbits below $\Gamma_{s}$ and “above” $W^{c}_{e}\cup\Gamma_{u}$ are asymptotic to $v_{\infty}$ as $t\to$-$\infty$. Orbits “below” $W^{c}_{e}\cup\Gamma_{u}$ are asymptotic to $r_{\infty}$ as $t\to$-$\infty$. There exists a critical value $\alpha_{c}\in(0.1,0.5)$ at which $W^{c}_{e}$ merges with $\Gamma_{s}$; at this bifurcation, the set of scattering orbits (present for $0\leq\alpha<\alpha_{c}$) disappears.
@@ 4.1.1 Phase portrait topology
To build the full phase portrait topology on the Poincaré disk, we considered three scenarios: $\alpha=0$, $\alpha=0.1$, and $\alpha=0.5$, all for $C=\gamma=1$ (see Fig. 4). In terms of the non-dimensional parameter of the system $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}$, equation (27), these parameters correspond respectively to: $c=\infty$, $c\approx 2.15$, and $c\approx 1.26$. We remark that the basin of attraction of $r_{c}$ increases along with $\alpha$, as expected.
@@ 4.2 The angular-momentum-energy space
An orbit in the conservative Kepler system maps to a single point in the CH space, as $C$ and $\mathcal{H}$ are constants. In the presence of dissipation, however, the image of an initial point, not on the curve of minimum energy, always moves vertically down. There are two possibilities for this motion as $t\to\infty$. One, where the point moves vertically down till it reaches the minimum energy curve, which happens if, and only if, the initial point is in the basis of attraction of the circular orbit. The second possibility is that the point moves vertically down until reaching a limit with $\mathcal{H}>0$. This scenario happens if, and only if, the initial point belongs either to the scattering region or to the escaping region.
4.3 Collisions and almost collisions.
In the dissipative Kepler problem considered here, collisions differ significantly from those in the conservative case.
Collisions occur only when the angular momentum CC is zero. In this case, independently of the value of the dissipative parameter $\alpha\geq 0$, the orbits lie on straight segments with the origin, which is the collision point, as an endpoint.
In the conservative case, an almost collision corresponds to an elliptic orbit with eccentricity $e$ close to one and one focus at the origin. Since $C=\sqrt{\gamma a(1-e^{2})}$, where $a$ is the semimajor axis, we have $C\approx 0$. The orbit then approximates a segment of length $2a$, with one endpoint (the periapsis) close to the origin. This suggests a collision regularization procedure: for $C=0$, the orbit becomes a segment of length $2a$ with the origin as one endpoint, and the particle moves periodically between the endpoints with infinite velocity at the periapsis.
In the following paragraphs we show that the dynamical behavior of collision orbits changes qualitatively when $\alpha>0$, regardless of how small $\alpha$ is.
This effect is a consequence of the specific structure of the dissipative force. The three parameters in equation (24) can be reduced to the single parameter $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}$, given in equation (27), by choosing $\alpha$ as the unit of time and $\gamma^{1/3}\alpha^{2/3}$ as the unit of length. With this scaling, system (24) becomes
$\begin{cases}\dot{r}=v\\ \dot{v}=\frac{c^{2}}{r^{3}}-\left(\frac{1}{r^{2}}+\frac{v}{r^{3}}\right)\end{cases}$ (42)
If $\alpha\ll 1$, the time $\tau$ in equation (42) is a slow time, related to the original time $t$ by $t=\alpha\tau$, while the length scale $\lambda$ corresponds to a magnified spatial variable $\ell=\gamma^{1/3}\alpha^{2/3}\lambda$. Velocities are correspondingly rescaled as $\frac{\Delta\ell}{\Delta t}=\frac{\gamma^{1/3}}{\alpha^{1/3}}\frac{\Delta\lambda}{\Delta\tau}$.
The key point is that, for $\alpha>0$, the limit $C\to 0$ (equivalently $c\to 0$) in equation (42) produces the same effective dynamics as the strong dissipative limit obtained by keeping $C>0$ fixed and letting $\alpha\to\infty$. Thus, in this model, close approaches between the bodies are intrinsically associated with strong energy dissipation.
We now return to the original time and length scales used in equation (24). For $\alpha>0$ and $C>0$ sufficiently small so that $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}$ is small, most solutions are attracted to the circular periodic orbit, as shown in Figure 4. In this regime, the radius $r_{c}=\frac{C^{2}}{\gamma}$ is small, and the motion rapidly settles into a circular orbit close to the origin.
For $c^{6}<\frac{1}{4}$, or $c\approx 0.79$, see equation (31), $(r_{c},0)$ is an attracting node. Thus, $r(t)$ with $r(0)>r_{c}$ converges to $r_{c}$ with at most one local maximum when $\dot{r}(0)>0$.
In contrast, for $\alpha=0$ and fixed small $C>0$, the motion is elliptical with eccentricity close to one. If $\alpha>0$ is introduced while keeping $c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}\gg 1$, the ellipse slowly shrinks toward the circular orbit of radius $r_{c}=\frac{C^{2}}{\gamma}$. Increasing $\alpha$ further so that $c<0.79$, or $\alpha>\frac{2C^{3}}{\gamma^{2}}$, leads to rapid convergence: most of the orbital energy is dissipated near the first perihelion passage, and $r(t)\to r_{c}$ with at most one oscillation. Note that as $C\to 0$, the threshold $\alpha>\frac{2C^{3}}{\gamma^{2}}$ also tends to zero.
At the same time, as $C\to 0$, the ejection equilibrium $(0,v_{e})=\left(0,\frac{C^{2}}{\gamma\alpha}\right)$ given in equation (30) converges to $(0,0)$. Hence, as $C\to 0$, $(0,v_{e})$ approaches $(r_{c},0)$, and these equilibria merge at $C=0$.
We now analyze the collision orbits, $C=0$, for any fixed value of $\alpha>0$. In this case, equation (42) becomes
$\begin{cases}\dot{r}=v\\ \dot{v}=-\left(\frac{1}{r^{2}}+\frac{v}{r^{3}}\right)\end{cases}$ (43)
As in equation (28), we regularize the vector field in equation (43) by multiplying it by $r^{3}$:
$\begin{cases}\dot{r}=r^{3}v\\ \dot{v}=-\left(r+v\right).\end{cases}$ (44)
Equation (44) has a single equilibrium $(r,v)=(0,0)$, with eigenvalues 0 and 1-1. This equilibrium admits an attracting center manifold tangent to the line v=rv=-r and a stable manifold tangent to the $v$-axis. A computation shows that the center manifold is given by
$v=h(r)=-r-r^{4}(1+5r^{3}+44r^{6}+539r^{9}+8337r^{12}+154632r^{15}\ldots)$ (45)
Since $h(r)<0$ for $r>0$, equation (44) implies that on the center manifold $\dot{r}=r^{3}h(r)<0$, so $(r,v)=(0,0)$, which represents collisions, is attracting.
The transverse attraction to the center manifold is stronger than the attraction along it; thus, solutions rapidly approach the center manifold and then drift along it toward $r=0$. Figure 6 shows the approximation of the center manifold (green) together with representative trajectories of (44).
Since the graph $v=h(r)$ is also invariant under the flow of equation (43), in the dissipative case $\alpha>0$ (with $\alpha$ normalized to one and length scale $\gamma^{1/3}\alpha^{2/3}$ equal to one), collisions occur with finite speed approximately given by equation (45), as $\dot{r}=v$.
Figure 6: The center manifold of $(r,v)=(0,0)$ as given in equation (45) and several stream lines of the vector field (44). The topology of the stream lines of equation (43) is the same.
In conclusion, while in the conservative case regularized collision orbits oscillate along a segment with infinite velocity at the origin, in the dissipative case, for any $\alpha>0$, a collision orbit with $\dot{r}(0)<0$ moves monotonically toward the origin with finite velocity.
@@ 6 Conclusion
Collision orbits in the conservative and dissipative settings are fundamentally different. While in the conservative case regularized collision orbits oscillate along a segment with infinite velocity at the origin, in the dissipative case, for any $\alpha>0$, a collision orbit with $\dot{r}(0)<0$ moves monotonically toward the origin with finite velocity.
As mentioned above, the Kepler system corresponds to the homographic dynamics of any planar central configuration under dissipation. For sufficiently small dissipation and nonzero angular momentum, configurations in the basins of attraction of the circular equilibrium have their homothetic (radial) motion damped, leading to purely rotational motion and hence to a relative equilibrium. In the dissipative case, purely homothetic solutions lead to collisions with finite velocity, as described above and in more detail in Section 4.3.
To extend the current model, one could consider the effect of rotation of bodies, reintroducing some of the neglected tidal effects. Moreover, to increase flexibility, one may also assign different dissipation strengths amongst pairs of bodies, $\alpha\to\alpha_{ij}=\alpha_{i}\alpha_{j}$, adapting to realistic scenarios where the primary bodies respond differently to tides due to its inner structure.